OpenStudy (anonymous):

Use the Quadratic Formula to solve the equation. –x2 + 7x = 8

OpenStudy (anonymous):

a lot easier to solve \[x^2-7x+8=0\]

OpenStudy (anonymous):

\[-x^2 + 7x = 8 \] Pretty sure that's what you're trying to write. Move everything to the right hand side to eliminate the minus sign of the x^2 and leave it in the form where it is EQUAL to ZERO. \[x^2-7x+8=0\] Use the quadratic formula and sub in the corresponding values. \[x=\frac{ -b \pm \sqrt{b^2-4ac} }{ 2a }\] \[x=\frac{ 7 \pm \sqrt{49-32} }{ 2 }\] \[x=\frac{7 \pm \sqrt{17} }{ 2 }\]

OpenStudy (anonymous):

\[7/2+- \sqrt{34}/2 \]

OpenStudy (anonymous):

is that it?

OpenStudy (anonymous):

Where did you get \[\sqrt{34}\] from?

OpenStudy (anonymous):

Wait its square root of 17. Not 34.

OpenStudy (anonymous):

Good job mate. Good luck for future questions.

OpenStudy (anonymous):

Thanks for helping (: